# Hidden Bose-Einstein Singularities in Correlated Electron Systems

### v1

@@ II.3 Superconducting state
The above consideration for the normal state can be extended directly and continuously to the $s$-wave superconducting state characterized by the emergence of the anomalous Green’s function $F(\vec{k})\!=\!F(-\vec{k})$. Specifically, we can construct the $\Phi$ functional in the FLEX-S approximation so as to incorporate all the anomalous processes derivable from those of Eq. (3) based on the diagrams of Fig. 1 Kita11 ,
@@
Matrix $\underline{\chi}^{({\rm c})}(\vec{q}\,)$ represents mixing of the superconducting and charge fluctuations due to the emergence of $F$. Equation (9) is now replaced by
@@
where $\lambda$ is determined by the condition that the smallest eigenvalue of the matrix $\underline{1}+U\underline{\chi}^{({\rm c})}(\vec{q}\,)$ stay zero once the value has been reached. Although it should be removed when all the eigenvalues are positive, our numerical study given below in Sect. III presents an example that incorporating $\varDelta\Phi_{\lambda}$ is indispensable to obtaining physical (i.e. finite) solutions throughout the superconducting phase.
The contributions of Eq. (12) to the normal and anomalous self-energies can be calculated by $\Sigma(\vec{k})\!=\!\frac{\beta}{2}\frac{\delta\Phi}{\delta G(\vec{k})}$ and $\Delta(\vec{k})\!=\!\frac{\beta}{2}\frac{\delta\Phi}{\delta F(\vec{k})}$, yielding Kita11
@@
Here $U_{+}^{\rm eff}(\vec{q}\,)$ is defined by
$U_{+}^{\rm eff}(\vec{q}\,)\equiv U\biggl{[}\frac{U\chi_{+}(\vec{q}\,)}{1-U\chi_{+}(\vec{q}\,)}-\frac{1}{3}U\chi_{+}(\vec{q}\,)\biggr{]},$ (16a)
@@
$\Sigma_{\lambda}(\vec{k})=$ $\,\Sigma(\vec{k})+U\frac{N}{V}\left[\lambda_{\bar{G}}G(-\vec{k})+\lambda_{G}G(\vec{k})+\lambda_{F}F(\vec{k})\right],$ (17a)
$\Delta_{\lambda}(\vec{k})=$ $\,\Delta(\vec{k})+U\frac{N}{V}\left\{\tilde{\lambda}_{F}F(\vec{k})+\tilde{\lambda}_{G}[G(\vec{k})+G(-\vec{k})]\right\},$ (17b)
@@
$\lambda_{\bar{G}}\!\equiv\!2\lambda\bigl{\{}\bigl{[}1\!-\!U\chi_{\bar{G}G}(\vec{0})\bigr{]}\bigl{[}1\!+\!U\chi_{-}(\vec{0})\bigr{]}\!-\!2\bigl{[}U\chi_{\bar{G}F}(\vec{0})\bigr{]}^{2}\bigr{\}},$
for example. The Dyson-Gor’kov equation can be written in terms of these self-energies as
@@
with $D(\vec{k})\!\equiv\![G_{0}^{-1}(\vec{k})\!-\!\Sigma_{\lambda}(\vec{k})][-G_{0}^{-1}(-\vec{k})\!+\!\Sigma_{\lambda}(-\vec{k})]\!-\![\Delta_{\lambda}(\vec{k})]^{2}$. It has turned out numerically that $\lambda_{\bar{G}}\!=\!\tilde{\lambda}_{F}\!>\!0$ and $\lambda_{G}\!=\!\lambda_{F}\!=\!\tilde{\lambda}_{G}\!=\!0$ holds for $\lambda$ that meets $\det[\underline{1}+U\underline{\chi}^{({\rm c})}(\vec{q}\,)]=0$. Note that $\lambda_{\bar{G}}$ should reach $\lambda_{\rm n}$ of Eq. (10) at the superconducting transition temperature. This completes our formulation to study the superconducting phase continuously from the normal state based on the FLEX-S approximation.
@@ III Numerical Results
Self-consistent calculations of Dyson’s and the Dyson-Gor’kov equations have been performed numerically on the negative-$U$ Hubbard model at a low electron density under the condition of the constant density
@@
where $0_{+}$ denotes an infinitesimal positive constant; the technical details are given in Appendix. We have adopted the units of $12m\!=\!\hbar\!=\!k_{\rm F}\!=\!$ 1, where $m$ is the electron mass and $k_{\rm F}$ is the Fermi momentum. The momentum cutoff $k_{\rm c}\!=\!10$ has been introduced to describe a finite band width. Specifically, our $G_{0}^{-1}(\vec{k})$ can be written as $G_{0}^{-1}(\vec{k})\!=\!i\varepsilon_{n}\!-\!k^{2}\!+\!\mu$ with $100\!\leq\!k\!\leq\!$ 10, where $\mu$ is the chemical potential determined by Eq. (19). We have chosen the parameter $U$ as $UN(0)=-0.09$ in terms of the density of states per spin and volume at the Fermi energy $N(0)\!=\!\frac{1}{4\pi^{2}}$; note also $\frac{N}{V}\!=\!\frac{1}{3\pi^{2}}$ in the present units. The corresponding mean-field superconducting transition temperature $T_{{\rm c}0}$, which is evaluated by using only the first terms on the right-hand sides of Eqs. (15a) and (15b), has the value $T_{{\rm c}0}\!=\!0.1767$ that is one order of magnitude smaller than the noninteracting Fermi energy $\varepsilon_{\rm F}^{0}\!=\!1$. Thus, our system lies in the intermediate coupling regime.
@@
Figure 2 plots temperature dependences of the Lagrange multipliers $\lambda_{\rm n}$ and $\lambda_{\bar{G}}$ of Eqs. (10) and (17a), respectively, together with the first-order anomalous self-energy $\Delta_{1}$ defined by the first term on the right-hand side of Eq. (15b) in terms of fully self-consistent $F$ with correlations, which is to be compared with the mean-field energy gap $\Delta_{\rm MF}$. The crosses and squares denote numerical results, and smooth curves are obtained by interpolating them. Matching the curves of $\lambda_{\rm n}$ and $\lambda_{\bar{G}}$ yields an estimate of the superconducting transition temperature $T_{{\rm c}}\!=\!0.0654$ with correlations, which is far below the mean-field value $T_{{\rm c}0}\!=\!0.1767$. This mean-field calculation yields the ratio $\Delta_{\rm MF}(0)/T_{{\rm c}0}\!=\!1.635$ that is still close to the weak-coupling value $1.76$ Leggett06 , while the one with correlations evaluated by $\Delta_{1}(0)/T_{{\rm c}}$ is enhanced up to $2.68$. Note that all the quantities in Fig. 2, which originally have the dimension of energy, are of the same order as $|UN(0)|=0.09$, as may be expected naturally.
@@
Figure 3: (a) Real and (b) imaginary parts of the correlation self-energy $\Sigma_{\rm c}(k,i\varepsilon_{n})$ calculated at $T\!=\!0.08$, which is expressed as a sum $\Sigma_{\rm c}\!=\!\Sigma_{\rm c}^{\rm i}\!+\!\Sigma_{\rm c}^{\rm r}$ of the 1PI and 1PR contributions $\Sigma_{\rm c}^{\rm i}$ and $\Sigma_{\rm c}^{\rm r}$, respectively.
@@
One of the key features of Fig. 2 is that $\lambda_{\rm n}$ starts to develop well above $T_{\rm c}\!=\!0.0654$ at $T_{1{\rm PR}}\!=\!0.100$, which means that the singularity of $U\chi_{G\bar{G}}(0)\!=\!1$ is reached already in the normal state. This is a continuous transition with no apparent broken symmetries, which is characterized by the emergence of the 1PR structure in the self-energy. Figure 3 shows (a) real and (b) imaginary parts of the correlation self-energy $\Sigma_{{\rm c}}$, i.e., Eq. (10) without the $\frac{UN}{2V}$ term in Eq. (5), calculated at $T\!=\!0.08$, where $\Sigma_{{\rm c}}^{\rm i}$ and $\Sigma_{{\rm c}}^{\rm r}$ denote the one-particle-irreducible (1PI) and 1PR contributions, respectively. Thus, the emergence of $\Sigma_{{\rm c}}^{\rm r}$ brings a marked change in the structure of the self-energy, especially around the Fermi momentum at low Matsubara frequencies, as expected naturally from $\Sigma_{{\rm c}}^{\rm r}(\vec{k})\!\propto\!G(-\vec{k})$.
@@
The emergence of the 1PR structure in the self-energy necessarily accompanies a change in the single-particle density of states per volume
$D(\varepsilon)\!\equiv\!-\frac{2}{\pi V}\sum_{{\bf k}}{\rm Im}\,G(k,i\varepsilon_{n}\rightarrow\varepsilon\!+\!i0_{+}).$ (20)
Figure 5 plots the density of states over $-0.5\!\leq\!\varepsilon\!\leq 0.5$ at five different temperatures obtained by a numerical analytic continuation of the self-energy based on the 6-point Padé approximant VS77 . As seen there, the reduction in the density of states around the excitation threshold, called the pseudogap behavior, starts even from above $T_{1{\rm PR}}\!=\!0.100$ and become more and more conspicuous down through $T_{1{\rm PR}}$. The singular 1PR structure makes a reliable analytic continuation based on the 6-point Padé approximant impossible below $T\!=\!0.098$, which is an issue to be clarified in the future. The present mechanism for the pseudogap behavior may also be relevant to those observed in high-$T_{\rm c}$ superconductors, which still remains a matter of controversy TS99 ; NPK05 ; Matsuda17 . One of the possible candidates relevant to it in the present context is the electron-hole bubble $\chi_{GG}({\bf K}/2)$ at half the reciprocal lattice vector ${\bf K}$ corresponding to the antiferromagnetism. The scenario that $U\chi_{GG}({\bf K}/2)=1$ with $U>0$ is reached upon the emergence of the pseudogap without any explicit broken symmetries may be well worthy of giving detailed theoretical investigations.
It should be noted finally that the concept of the 1PR structure in the self-energy has emerged naturally in our effort of describing strongly correlated systems microscopically and numerically; without introducing Eqs. (9) and (14), we could not have obtained any finite solution of Dyson’s and the Dyson-Gor’kov equations below $T_{1{\rm PR}}$. In this respect, our model with a definite cutoff $k_{\rm c}\!=\!10$ may have been advantageous over the continuous one with ultraviolet divergences HRCZ07 in that no further renormalization of $U$, which may mask the singularities, is required.
IV Concluding Remarks
We have clarified existence of hidden singularities in correlated electron systems and developed a definite procedure to handle them. It is thereby shown that reaching them is characterized by the emergence of a novel structure in the self-energy called one-particle reducible (1PR), which necessarily accompanies a change in the single-particle density of states around the excitation threshold. The temperature at which it starts to develop can be regarded as a transition with no apparent broken symmetries.
The hidden singularities may have stood as a key untold difficulty in performing theoretical studies on correlated ordered phases based on the quantum field theory, as may be seen in the negligible number of theoretical studies on them compared with those on the normal state Yanase03 ; Kontani13 ; Scalapino12 , a resolution for which has been presented here. The 1PR structure has also been predicted to present in Bose-Einstein condensates based on a self-consistent $\Phi$ derivable scheme of satisfying Goldstone’s theorem and conservation laws Kita09 ; Kita14 . Thus, it may be a common feature over a wide range of correlated systems. Further detailed studies will be required to fully elucidate the novel phase with the 1PR structure in the self-energy.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant Number JP20K03848.
Appendix A Numerical Procedures

### v2

@@ II.3 Superconducting state
The above consideration for the normal state can be extended directly and continuously to the $s$-wave superconducting state characterized by the emergence of the anomalous Green’s function $F(\vec{k})\!=\!F(-\vec{k})$. Specifically, we can construct the $\Phi$ functional in the FLEX-S approximation so as to incorporate all the anomalous processes derivable from those of Eq. (3) based on the diagrammatic expression in Fig. 1 Kita11 ,
@@
The matrix $\underline{\chi}^{({\rm c})}(\vec{q}\,)$ represents the mixing of superconducting and charge fluctuations due to the emergence of $F$. Equation (9) is now replaced by
@@
where $\lambda$ is determined by the condition that the smallest eigenvalue of the matrix $\underline{1}+U\underline{\chi}^{({\rm c})}(\vec{q}\,)$ stays zero once this has been reached. Although it should be removed when all the eigenvalues are positive, our numerical study given below in Sect. III presents an example that incorporating $\varDelta\Phi_{\lambda}$ is indispensable to obtaining physical (i.e., finite) solutions throughout the superconducting phase.
The contributions of Eq. (12) to the normal and anomalous self-energies can be determined by calculating $\Sigma(\vec{k})\!=\!\frac{\beta}{2}\frac{\delta\Phi}{\delta G(\vec{k})}$ and $\Delta(\vec{k})\!=\!\frac{\beta}{2}\frac{\delta\Phi}{\delta F(\vec{k})}$, yielding Kita11
@@
Here, $U_{+}^{\rm eff}(\vec{q}\,)$ is defined by
$U_{+}^{\rm eff}(\vec{q}\,)\equiv U\biggl{[}\frac{U\chi_{+}(\vec{q}\,)}{1-U\chi_{+}(\vec{q}\,)}-\frac{1}{3}U\chi_{+}(\vec{q}\,)\biggr{]}$ (16a)
@@
$\Sigma_{\lambda}(\vec{k})=$ $\,\Sigma(\vec{k})+\bar{n}U\!\left[\lambda_{\bar{G}}G(-\vec{k})+\lambda_{G}G(\vec{k})+\lambda_{F}F(\vec{k})\right],$ (17a)
$\Delta_{\lambda}(\vec{k})=$ $\,\Delta(\vec{k})+\bar{n}U\!\left\{\tilde{\lambda}_{F}F(\vec{k})+\tilde{\lambda}_{G}[G(\vec{k})+G(-\vec{k})]\right\},$ (17b)
@@
$\lambda_{\bar{G}}\!\equiv$  $\,2\lambda\bigl{\{}\bigl{[}1\!+\!U\chi_{-}(\vec{0})\bigr{]}\bigl{[}1\!-\!U\chi_{G\bar{G}}(\vec{0})\bigr{]}\!-\!2\bigl{[}U\chi_{GF}(\vec{0})\bigr{]}^{2}\bigr{\}},$
$\lambda_{G}\!\equiv$ $\,-\lambda\bigl{\{}\bigl{[}1\!-\!U\chi_{G\bar{G}}(\vec{0})\bigr{]}^{2}\!-\!\bigl{[}U\chi_{FF}(\vec{0})\bigr{]}^{2}\bigr{\}},$
$\lambda_{F}\!\equiv$  $\,4\lambda\,U\chi_{GF}(\vec{0})\bigl{[}1\!-\!U\chi_{G\bar{G}}(\vec{0})\!-\!U\chi_{FF}(\vec{0})\bigr{]},$
$\tilde{\lambda}_{F}\!\equiv$  $\,2\lambda\bigl{\{}\bigl{[}1\!+\!U\chi_{-}(\vec{0})\bigr{]}U\chi_{FF}(\vec{0})\!-\!2\bigl{[}U\chi_{GF}(\vec{0})\bigr{]}^{2}\bigr{\}}$
$\,+\lambda\bigl{\{}\bigl{[}1\!-\!U\chi_{G\bar{G}}(\vec{0})\bigr{]}^{2}\!-\!\bigl{[}U\chi_{FF}(\vec{0})\bigr{]}^{2}\bigr{\}},$
$\tilde{\lambda}_{G}\!\equiv$  $\,2\lambda\,U\chi_{GF}(\vec{0})\bigl{[}1\!-\!U\chi_{G\bar{G}}(\vec{0})\!-\!U\chi_{FF}(\vec{0})\bigr{]}.$
The Dyson-Gor’kov equation can be written in terms of these self-energies as
@@
with $D(\vec{k})\!\equiv\![G_{0}^{-1}(\vec{k})\!-\!\Sigma_{\lambda}(\vec{k})][-G_{0}^{-1}(-\vec{k})\!+\!\Sigma_{\lambda}(-\vec{k})]\!-\![\Delta_{\lambda}(\vec{k})]^{2}$. It has turned out numerically that $\lambda_{\bar{G}}\!=\!\tilde{\lambda}_{F}\!>\!0$ and $\lambda_{G}\!=\!\lambda_{F}\!=\!\tilde{\lambda}_{G}\!=\!0$ hold for $\lambda$ that meets $\det[\underline{1}+U\underline{\chi}^{({\rm c})}(\vec{q}\,)]=0$. Note that $\lambda_{\bar{G}}$ should reach $\lambda_{\rm n}$ of Eq. (10) at the superconducting transition temperature. This completes our formulation to study the superconducting phase continuously from the normal state on the basis of the FLEX-S approximation.
@@ III Numerical Results
Self-consistent calculations of the Dyson and Dyson-Gor’kov equations have been performed numerically on the negative-$U$ Hubbard model at a low electron density under the constant density
@@
where $0_{+}$ denotes an infinitesimal positive constant; the technical details are given in AppendixB. We have adopted the units of $12m\!=\!\hbar\!=\!k_{\rm F}\!=\!$ 1, where $m$ is the electron mass and $k_{\rm F}$ is the Fermi momentum. The momentum cutoff $k_{\rm c}\!=\!10$ has been introduced to describe a finite band width. Specifically, our $G_{0}^{-1}(\vec{k})$ can be written as $G_{0}^{-1}(\vec{k})\!=\!i\varepsilon_{n}\!-\!k^{2}\!+\!\mu$ with $100\!\leq\!k\!\leq\!$ 10, where $\mu$ is the chemical potential determined by Eq. (19). We have chosen the parameter $U$ as $UN(0)=-0.09$ in terms of the density of states per spin and the volume at the Fermi energy $N(0)\!=\!\frac{1}{4\pi^{2}}$; note also that $\frac{N}{V}\!=\!\frac{1}{3\pi^{2}}$ in the present units. The corresponding mean-field superconducting transition temperature $T_{{\rm c}0}$, which is evaluated by using only the first terms on the right-hand sides of Eqs. (15a) and (15b), is $0.1767$ that is one order of magnitude smaller than the noninteracting Fermi energy $\varepsilon_{\rm F}^{0}\!=\!1$. Thus, our system lies in the intermediate coupling regime.
@@
Figure 2 shows the temperature dependences of the Lagrange multipliers $\lambda_{\rm n}$ and $\lambda_{\bar{G}}$ of Eqs. (10) and (17a), respectively, together with the first-order anomalous self-energy $\Delta_{1}$ defined by the first term on the right-hand side of Eq. (15b) in terms of the fully self-consistent $F$ with correlations, which is to be compared with the mean-field energy gap $\Delta_{\rm MF}$. The crosses and squares denote numerical results, and smooth curves are obtained by interpolating them. Matching the curves of $\lambda_{\rm n}$ and $\lambda_{\bar{G}}$ yields an estimate of the superconducting transition temperature $T_{{\rm c}}\!=\!0.0654$ with correlations, which is far below the mean-field value $T_{{\rm c}0}\!=\!0.1767$. This mean-field calculation yields the ratio $\Delta_{\rm MF}(0)/T_{{\rm c}0}\!=\!1.635$ that is still close to the weak-coupling value of $1.76$ Leggett06 , whereas that with correlations evaluated by $\Delta_{1}(0)/T_{{\rm c}}$ is increased up to $2.68$. Note that all the quantities in Fig. 2, which originally have the dimension of energy, are of the same order as $|UN(0)|=0.09$, as may be expected naturally.
@@
Figure 3: (a) Real and (b) imaginary parts of the correlation self-energy $\Sigma_{\rm c}(k,i\varepsilon_{n})$ calculated at $T\!=\!0.08$, which is expressed as the sum $\Sigma_{\rm c}\!=\!\Sigma_{\rm c}^{\rm i}\!+\!\Sigma_{\rm c}^{\rm r}$ of the 1PI and 1PR contributions $\Sigma_{\rm c}^{\rm i}$ and $\Sigma_{\rm c}^{\rm r}$, respectively.
@@
One of the key features of Fig. 2 is that $\lambda_{\rm n}$ starts to develop well above $T_{\rm c}\!=\!0.0654$ at $T_{1{\rm PR}}\!=\!0.100$, which means that the singularity of $U\chi_{G\bar{G}}(0)\!=\!1$ is reached already in the normal state. This is a continuous transition with no apparent broken symmetries, which is characterized by the emergence of the 1PR structure in the self-energy. Figure 3 shows the (a) real and (b) imaginary parts of the correlation self-energy $\Sigma_{{\rm c}}$, i.e., Eq. (10) without the $\frac{UN}{2V}$ term in Eq. (5), calculated at $T\!=\!0.08$, where $\Sigma_{{\rm c}}^{\rm i}$ and $\Sigma_{{\rm c}}^{\rm r}$ denote the one-particle-irreducible (1PI) and 1PR contributions, respectively. Thus, the emergence of $\Sigma_{{\rm c}}^{\rm r}$ brings a marked change in the structure of the self-energy, especially around the Fermi momentum at low Matsubara frequencies, as expected naturally from $\Sigma_{{\rm c}}^{\rm r}(\vec{k})\!\propto\!G(-\vec{k})$.
@@
The emergence of the 1PR structure in the self-energy is necessarily accompanied by a change in the single-particle density of states per volume,
$D(\varepsilon)\!\equiv\!-\frac{2}{\pi V}\sum_{{\bm{k}}}{\rm Im}\,G(k,i\varepsilon_{n}\rightarrow\varepsilon\!+\!i0_{+}).$ (20)
Figure 5 shows the density of states over $-0.5\!\leq\!\varepsilon\!\leq 0.5$ at five different temperatures obtained by a numerical analytic continuation of the self-energy based on the six-point Padé approximant VS77 . As seen in this figure, the reduction in the density of states around the excitation threshold, called the pseudogap behavior, starts even from above $T_{1{\rm PR}}\!=\!0.100$ and becomes increasingly conspicuous down through $T_{1{\rm PR}}$. The singular 1PR structure makes a reliable analytic continuation based on the six-point Padé approximant impossible below $T\!=\!0.098$, which is an issue to be clarified in the future. The present mechanism for the pseudogap behavior may explain those observed in ultracold atomic gases.SGJ08 ; GSDJPPS10 In this respect, it is worth noting that the pseudogap behavior has been reproduced qualitatively by numerical calculations on the negative-$U$ Hubbard model based on the dynamical mean-field theory,PB15 ; Sakai15 and attributed to the emergence of a pole in the self-energy in the strong-coupling low-temperature region;Sakai15 the present study can trace the microscopic origin of this pole to the second term of Eq. (10). Since the hidden singularity that we have found is general and may not depend on microscopic details of the model, it is interesting to explore in the future whether the present mechanism can explain the pseudogap observed in cuprate superconductorsTS99 ; NPK05 ; Matsuda17 by taking an appropriate ${\bm{q}}$ for either $U\chi_{G\bar{G}}({\bm{q}},0)\!=\!1$ or $U\chi_{GG}({\bm{q}},0)\!=\!1$.
@@
Note finally that the concept of the 1PR structure in the self-energy has emerged naturally in our effort to describe strongly correlated systems microscopically and numerically; without introducing Eqs. (9) and (14), we could not have obtained any finite solution of the Dyson and Dyson-Gor’kov equations below $T_{1{\rm PR}}$. In this respect, our model with a definite cutoff $k_{\rm c}\!=\!10$ may have been advantageous over the continuous model with ultraviolet divergences HRCZ07 in that no further renormalization of $U$, which may mask the singularities, is required.
IV Conclusions
We have clarified existence of hidden singularities in correlated electron systems and developed a definite procedure to handle them. It is thereby shown that reaching the singularities is characterized by the emergence of a novel structure in the self-energy called 1PR, which is necessarily accompanied by a change in the single-particle density of states around the excitation threshold. The temperature at which it starts to develop can be regarded as a transition with no apparent broken symmetries.
They singularities are inherent in the interaction channel and should also present in Bose systems. They may have stood as a key untold difficulty in performing theoretical studies on correlated normal and ordered phases based on the quantum field theory, as may be seen in the negligible number of theoretical studies on them compared with those on the normal state Yanase03 ; Kontani13 ; Scalapino12 , a resolution for which has been presented here. The 1PR structure has also been predicted to be present in Bose-Einstein condensates on the basis of a self-consistent $\Phi$ derivable scheme of satisfying Goldstone’s theorem and conservation laws Kita09 ; Kita14 . Thus, it may be a common feature over a wide range of correlated systems. Further detailed studies will be required to fully elucidate the novel phase with the 1PR structure in the self-energy.
@@ Acknowledgements.
The author is grateful to Masatoshi Imada for his valuable comments on the present results. This work was supported by JSPS KAKENHI Grant Number JP20K03848.
Appendix A General Definition of Hidden Singularities
Figure 6: Diagrammatic expression of $\Gamma^{(\rm ir)}$ defined by Eq. (21). The left (right) pair of external lines represents $11^{\prime}_{ii^{\prime}}$ ($22^{\prime}_{jj^{\prime}}$).
Let us define the irreducible vertex by Eq. (39) of Ref. Kita11, as
$\Gamma_{ii^{\prime},jj^{\prime}}^{(\rm ir)}(11^{\prime},22^{\prime})\equiv 2\beta\frac{\delta^{2}\Phi}{\delta G_{i^{\prime}i}(1^{\prime},1)\delta G_{j^{\prime}j}(2^{\prime},2)},$ (21)
where $1$ is defined by $1\!\equiv\!(\xi_{1},\tau_{1})$ with $\xi_{1}$ denoting the space-spin coordinates and $\tau_{1}\!\in\![0,\beta]$, and the subscript $i\!=\!1,2$ distinguishes the creation and annihilation operators. The differentiation of Eq. (21) amounts to cutting a pair of lines in Fig. 1 in all possible ways. Hence, the irreducible vertex can be expressed diagrammatically as shown in Fig. 6, which has the structure called simpleDDM64-2 or two-particle irreducible with respect to $11^{\prime}_{ii^{\prime}}$ and $22^{\prime}_{jj^{\prime}}$; see the paragraph below Eq. (43) in Ref. DDM64-2, . We also introduce the matrices $\underline{\Gamma}^{(\rm ir)}$, $\underline{\chi}$, and $\underline{1}$ as
$\langle 11^{\prime}_{ii^{\prime}}|\underline{\Gamma}^{(\rm ir)}|22^{\prime}_{jj^{\prime}}\rangle\equiv$ $\,{\Gamma}^{(\rm ir)}_{ii^{\prime},j^{\prime}j}(11^{\prime},2^{\prime}2),$ (22a)
$\langle 11^{\prime}_{ii^{\prime}}|\underline{\chi}|22^{\prime}_{jj^{\prime}}\rangle\equiv$ $\,-G_{ij}(1,2)G_{j^{\prime}i^{\prime}}(2^{\prime},1^{\prime}),$ (22b)
$\langle 11^{\prime}_{ii^{\prime}}|\underline{1}|22^{\prime}_{jj^{\prime}}\rangle\equiv$ $\,\delta_{ij}\delta_{i^{\prime}j^{\prime}}\delta(1,2)\delta(1^{\prime},2^{\prime}),$ (22c)
in the same way as Eqs. (25) and (42d) in Ref. Kita11, . Using them, we can generalize Eq. (28) in Ref. Kita11, in the FLEX-S approximation to obtain the whole logarithmic contribution to $\Phi$:
$\Phi_{\rm log}=\frac{1}{2\beta}{\rm Tr}\ln\bigl{(}\underline{1}+\underline{\Gamma}^{(\rm ir)}\underline{\chi}\bigr{)}.$ (23)
We then transform the $\tau$ variables in the brackets of Eq. (22a) into Matsubara frequencies as
$\langle 11^{\prime}|\underline{\Gamma}^{(\rm ir)}|22^{\prime}\rangle=$ $\,\frac{1}{\beta^{3}}\sum_{\ell}\sum_{n_{1}n_{2}}\langle\tilde{1}\tilde{1}^{\prime}_{ii^{\prime}}|\underline{\tilde{\Gamma}}^{(\rm ir)}(\omega_{\ell})|\tilde{2}\tilde{2}^{\prime}_{jj^{\prime}}\rangle$
×$\,\times e^{-i\varepsilon_{1}\tau_{1}+i\varepsilon_{1-}\tau_{1^{\prime}}+i\varepsilon_{2}\tau_{2}-i\varepsilon_{2-}\tau_{2^{\prime}}},$ (24)
where $\underline{\tilde{\Gamma}}^{(\rm ir)}$ denotes a matrix with the index $|\tilde{1}\tilde{1}^{\prime}_{ii^{\prime}}\rangle\equiv|\xi_{1}\xi_{1^{\prime}},\varepsilon_{n_{1}},ii^{\prime}\rangle$, the symbol $\varepsilon_{1}$ is the abbreviation for $\varepsilon_{n_{1}}$, and $\varepsilon_{1-}$ is defined by $\varepsilon_{1-}\equiv\varepsilon_{n_{1}}-\omega_{\ell}$. The matrices of Eqs. (22b) and (22c) can be expanded similarly in terms of the coefficients
$\langle\tilde{1}\tilde{1}^{\prime}_{ii^{\prime}}|\underline{\tilde{\chi}}(\omega_{\ell})|\tilde{2}\tilde{2}^{\prime}_{jj^{\prime}}\rangle\equiv$ $\,-\beta\delta_{n_{1}n_{2}}G_{ij}(\xi_{1},\xi_{2};\varepsilon_{1})$
×$\,\times G_{j^{\prime}i^{\prime}}(\xi_{2^{\prime}},\xi_{1^{\prime}};\varepsilon_{n_{1}}\!-\!\omega_{\ell}),$ (25a)
$\langle\tilde{1}\tilde{1}^{\prime}_{ii^{\prime}}|\underline{\tilde{1}}|\tilde{2}\tilde{2}^{\prime}_{jj^{\prime}}\rangle\equiv$ $\,\beta\delta_{n_{1}n_{2}}\delta_{ij}\delta_{i^{\prime}j^{\prime}}\delta(\xi_{1},\xi_{2})\delta(\xi_{1^{\prime}},\xi_{2^{\prime}}).$ (25b)
Using the tilde matrices, we can express Eq. (23) as
$\Phi_{\rm log}=$ $\,\frac{1}{2\beta}\sum_{\ell}{\rm Tr}\,\ln\Bigl{[}\underline{\tilde{1}}+\underline{\tilde{\Gamma}}^{(\rm ir)}(\omega_{\ell})\,\underline{\tilde{\chi}}(\omega_{\ell})\Bigr{]}$
$=$ $\,\frac{-1}{2\beta}\sum_{\ell}{\rm Tr}\,\ln\Bigl{[}\underline{\tilde{1}}-\underline{\tilde{\Gamma}}(\omega_{\ell})\,\underline{\tilde{\chi}}(\omega_{\ell})\Bigr{]},$ (26)
where $\underline{\tilde{\Gamma}}$ is the full vertex that obeys $\underline{\tilde{\Gamma}}\!=\!\underline{\tilde{\Gamma}}^{(\rm ir)}-\underline{\tilde{\Gamma}}^{(\rm ir)}\underline{\tilde{\chi}}\,\underline{\tilde{\Gamma}}$. The hidden singularity is defined generally as the point where the determinant of $\underline{\tilde{1}}+\underline{\tilde{\Gamma}}^{(\rm ir)}(0)\,\underline{\tilde{\chi}}(0)$ reaches zero, or equivalently, the determinant of $\underline{\tilde{1}}-\underline{\tilde{\Gamma}}(0)\,\underline{\tilde{\chi}}(0)$ diverges. Reducing the matrix by removing the singularity, one may find multiple singularities.
By either cutting the series in Fig. 6 at some finite order or taking its partial sum to infinity, we obtain an approximate treatment of the logarithmic singularity of Eq. (26), such as the FLEX-S approximation of Eqs. (3) and (12) that cuts the series in Fig. 6 at the lowest order. Alternatively, the parameter $U$ in Eqs. (3) and (12) may be regarded as an effective potential for the full irreducible vertex of Eq. (26).
Note finally that the hidden singularities do not specify second-order transition points with broken symmetries. For example, the superconducting transition temperature of the isotropic $s$-wave pairing is determined by Eq. (6) in terms of $U^{\rm s}_{\vec{k}\vec{k}^{\prime}}$ defined generally by $U^{\rm s}_{\vec{k}\vec{k}^{\prime}}\!\equiv\!\frac{\beta}{2}\frac{\delta^{2}\Phi}{\delta F(\vec{k})\delta F(\vec{k}^{\prime})}\Bigr{|}_{F=0}$.
Appendix B Numerical Procedures